### v1

@@ Abstract
Software bloat refers to code and features that is not used by a software during runtime. For Machine Learning (ML) systems, bloat is a major contributor to their technical debt leading to decreased performance and resource wastage.
In this work, we present, Negativa-ML, a novel tool to identify and remove bloat in ML frameworks by analyzing their shared libraries. Our approach includes novel techniques to detect and locate unnecessary code within device code - a key area overlooked by existing research, which focuses primarily on host code. We evaluate Negativa-ML using four popular ML frameworks across ten workloads over 300 shared libraries. The results demonstrate that the ML frameworks are highly bloated on both the device and host code side. On average, Negativa-ML reduces the device code size in these frameworks by up to 75% and the host code by up to 72%, resulting in total file size reductions of up to 55%. The device code is a primary source of bloat within ML frameworks. Through debloating, we achieve reductions in peak host memory usage, peak GPU memory usage, and execution time by up to 74.6%, 69.6%, and 44.6%, respectively.
@@ 1 Introduction
Machine learning (ML) has affected many industries significantly. From personalized recommendations Ko et al. (2022), to healthcare diagnostics Bhardwaj et al. (2017), and autonomous vehicles Parekh et al. (2022), ML is revolutionizing nearly every sector. Not only Large Language Models (LLMs) that enabled natural language understanding and content generation like ChatGPT OpenAI et al. (2024) and Llama Touvron et al. (2023b) have propelled this transformation, but also smaller models that are now widely deployed in many vision applications from cameras to mobile phones. However, as these models grow in numbers, size, and complexity, managing and optimizing ML systems has become increasingly challenging. LLMs often contain billions of parameters, requiring substantial computational resources and vast amounts of training data. Some smaller models are now deployed on embedded devices. To manage this complexity, substantial technical overhead is introduced to ML systems, exacerbating the issue of software bloat within these systems Zhang et al. (2024).
@@
Software bloat refers to code that is unnecessary for a program during runtime, typically caused by extraneous functions, libraries, or features that do not contribute to the core functionality. Software bloat can cause a range of issues, including decreased performance, increased resource usage, and security vulnerabilities. While bloat can affect any type of software, ML systems have a special capacity for incurring it, because they have all the maintenance problems of traditional code plus an additional set of ML-specific issues, such as boundary erosion, data dependencies, etc. Sculley et al. (2015). As ML models grow in scale and complexity, the bloat in ML systems compounds, leading to further inefficiencies that hinder runtime performance and increase the cost of deployment, particularly in resource-constrained environments Zhang et al. (2024); Jiang et al. (2020).
@@
At the heart of ML systems are ML frameworks, such as TensorFlow Abadi et al. (2016) and PyTorch Paszke et al. (2019). These frameworks provide the essential libraries and tools for model training and inference. ML frameworks are typically developed using multiple programming languages with C++ and CUDA used to implement the core functionalities, focusing on maximizing performance. Both C++ and CUDA code are compiled into shared libraries. Python acts as the frontend, wrapping these core functionalities and enhancing the frameworks’ usability for developers. However, with the ease of use and flexibility that these frameworks provide, they also introduce framework tax Fernandez et al. (2023) - these frameworks can introduce performance overheads and diminish the benefits of new hardware and model architecture advancements. Furthermore, integrating these frameworks also leads to large binary sizes and brings unnecessary overhead for mobile and edge devices Jiang et al. (2020).
@@
In this paper, we aim to identify and reduce bloat in ML frameworks, targeting the bloat in ML shared libraries. These shared libraries encapsulate the core functionalities of ML frameworks and constitute the majority of ML framework size. Debloating ML shared libraries involves the following challenges:
@@
  • ML frameworks rely on some proprietary libraries, such as cuDNN nvi (b) and cuBLAS nvi (c). These libraries are not open-sourced, making it impossible to perform source code analysis.
@@
  • ML frameworks contain code that runs on both the host (CPU) and device (GPU). Device code is overlooked by existing research and no method measure and reduce bloat in it.
@@
  • Although host code has a well-defined structure and the bloat in it has been previously studied anonymous , the structure of device code is not publicly available, making it difficult to analyze and debloat.
@@
To address these challenges, we propose a tool, Negativa-ML, to identify and remove bloat in both host and device code within ML shared libraries. Leveraging insights into how ML systems execute ML workloads on host and device, we propose a novel approach to detect the device code used by an ML workload with low performance overhead. Subsequently, we locate the file ranges occupied by this device code within ML shared libraries. This involves a deep understanding of how device code is compiled and how they are organized within a shared library, which is challenging because device code structure lacks a public specification. Finally, we utilize a debloating tool that we have previously developed anonymous to remove bloat from ML shared libraries according to the file ranges located. Our contributions are as follows:
@@
  • We propose an approach to detect device code used by ML workloads with low performance overhead.
@@
  • We analyze the structure of device code in ML shared libraries and propose a method to locate the file ranges of used device code.
@@
  • We extend a debloating tool that we have developed anonymous to remove bloat in both host and device code in ML shared libraries.
@@
To the best of our knowledge, this is the first work to evaluate and remove bloat in device code. Our evaluation shows that ML frameworks are highly bloated, with 72% of host code and 75% of device code being unnecessary for target ML workloads. Besides, 10% shared libraries account for up to 90% of the total bloat in the ML frameworks. This bloat not only increases storage overhead but also degrades runtime performance with increased host memory usage, GPU memory usage, and longer execution times.
@@ 2.1 ML Frameworks
ML frameworks are software libraries that provide essential tools for building, training, and deploying ML models. Some ML frameworks are general-purpose and designed to support both training and inference for a wide variety of ML models. While others are optimized for specific use cases, such as LLM inference. Two of the most popular general-purpose ML frameworks are TensorFlow Abadi et al. (2016) and PyTorch Paszke et al. (2019), both of which are widely adopted in industry and academia. Although the two frameworks target various ML models,
the rise of LLMs has introduced new requirements for ML frameworks, such as efficient KV cache management Kwon et al. (2023); Sinha et al. (2024). LLM inference — which is highly latency-sensitive and resource-demanding — is particularly challenging. To meet these demands, new frameworks specifically designed for LLM inference have been proposed Contributors (2023); Kwon et al. (2023); Wolf et al. (2020); hug .
@@
The standard file format for shared libraries is the Executable and Linkable Format (ELF) ora , which organizes a file into various sections, such as .text and .data. A generic shared library only contains code running on CPU in the .text section. This CPU-running code is called host code. However, ML shared libraries are unique in that they also contain the code designed to run on GPUs, which is referred to as device code. This device code is usually included in another section called .nv_fatbin in an ML shared library.
Device code significantly increases the size of ML shared libraries compared to traditional ones. The core shared library in PyTorch, torch.so, is 881 MB for the GPU version and 482 MB for the CPU version, nearly double the size. Furthermore, device code makes up a substantial portion of these libraries. Analyzing the top four largest shared libraries in PyTorch, device code accounts for between 68% to more than 91% of the size of each shared library, as shown in Figure 1. In contrast, host code represents only a small fraction of these libraries. Despite device code forming the majority of ML shared libraries, no existing research has investigated bloat within device code.
Figure 1: Distribution of host code and device code in the top 4 largest shared libraries in PyTorch.
@@ 2.2 Software Bloat
Software bloat is mainly caused by unnecessary code in a software, which can be categorized into two types Brown et al. (2024): Type I bloat is universally unnecessary, for example, dead and unreachable code. Type II bloat is end-use dependent, code may be or may not be Type II bloat depending on the use case, for example, unnecessary program features. Software bloat exists across all the modern software stack, from operating systems Quach et al. (2017); Kuo et al. (2020), shared libraries Ziegler et al. (2019); Quach et al. (2018); Biswas et al. (2021), to executables Navas & Gehani (2023); Ahmad et al. (2021); Qian et al. (2019) and even containerized applications Rastogi et al. (2017); Zhang et al. (2024). Bloat causes software to grow in size and complexity over time, which leads to performance degradation, increased memory consumption, longer startup time, and security vulnerabilities, but without providing any additional benefits.
Recently, bloat in ML systems has attracted increasing attention. Sculley et al. Sculley et al. (2015) identify that ML systems are easy to accumulate hidden technical debt due to a set of ML-specific issues. Zhang et al. Zhang et al. (2024) show that containerized ML applications are significantly bloated, wasting storage resources and network bandwidth, increasing their attack surface, and slowing down their deployment process.
@@ 2.3 Software Debloating
Software debloating is the process of removing unnecessary code from software to improve efficiency and reduce resource consumption. Many debloating tools have been developed for traditional software, which can be categorized by their debloating targets: source code, binary code, and containerized applications. Source code debloating tools eliminate unused code directly from the source, such as dead or unreachable code, based on specific usage scenarios Brown & Pande (2019); Ye et al. (2021); Azad et al. (2019). Binary debloating tools operate on software binaries, including shared libraries and executables, to remove unnecessary functions or instructions Qian et al. (2019); Ahmad et al. (2021); Ziegler et al. (2019); Agadakos et al. (2019). Container debloating tools target containerized applications, removing unnecessary files in container images Rastogi et al. (2017); Zhang et al. (2023). Most of these tools remove code that is not used by the specific workload, i.e., Type II bloat Qian et al. (2019); Ahmad et al. (2021); Brown & Pande (2019); Rastogi et al. (2017); Azad et al. (2019). While debloating traditional software has been extensively studied, to the best of our knowledge, all existing work focus only on traditional applications where code runs only on CPUs. Device code, which constitutes a significant portion of ML frameworks, remains unexamined.
One major issue with debloating tools that focus on binary debloating is that they are generally unreliable Brown et al. (2024). To solve their reliability issues, we have recently developed Negativa, a debloating tools that only debloat host code anonymous . [^1] Please see the supplementary material for details of this tool. Negativa demonstrates effectiveness in debloating shared libraries, profiling workloads with the target shared libraries, then removing any code not used by the workload. Negativa’s debloating process involves three phases: detection, location, and compaction. In the detection phase, it identifies host functions used by the workload; in the location phase, it locates the used host functions within the shared library; and in the compaction phase, it removes unused functions from the shared library and only keeps the used ones.
In this work, we introduce Negativa-ML, a technique that we have developed within Negativa to enable device code debloating. In addition, we build on Negativa’s capabilities to also debloat host code in ML shared libraries. Building on Negativa’s three-phase debloating process, we introduce novel approaches for detecting and locating used device code in the detection and location phases. Negativa’s compaction phase is then reused to remove unused device code.
3 Convention
For the clarity and consistency of the paper, we define the following terms: Host refers to CPU and its memory; Host code refers to the code that runs on the host; Functions specifically refer to the functions in host code. Device refers to NVIDIA GPUs running on CUDA drivers and its memory; Device code refers to the code that runs on the device; Kernels refer to the functions in device code.
@@ 4 Methodology
Figure 2 provides an overview of Negativa-ML. Negativa-ML is composed of two components: the kernel detector and the kernel locator. During the execution of an ML workload, such as model training with PyTorch, many shared librarfies of ML frameworks are utilized. The kernel detector monitors the kernel execution of the workload and records the names of the kernels used. The kernel locator analyzes the shared libraries, extracting device code within these libraries and finding the file ranges of the used kernels within the shared library. These ranges are subsequently passed to Negativa’s compaction module for debloating. Finally, a debloated ML shared library with reduced host code and device code is generated.
@@ 4.1 Kernel Detector
The kernel detector is responsible for detecting the kernels used by the ML workload. While existing tools like Nsight Systems (NSys) nvi (f) can profile device code, they are designed primarily for profiling and debugging rather than for used kernel detection. These tools, to provide comprehensive information, usually impose a high performance overhead on the target applications by recording data each time a kernel is called. However, for kernel detection, we are only interested in determining whether a kernel has been used, without needing repetitive call data, which results in unnecessary overhead. To offer a more efficient solution, we propose a novel, lightweight approach that specifically detects kernels used within ML shared libraries, achieving low performance overhead. Since kernels execute on GPU devices, an intuitive approach would be to monitor GPU device execution directly. However, this approach is not feasible because the code execution on GPU devices is not directly accessible.
@@
Our approach leverages insights into how ML systems execute ML workloads on both the host and the device; Considering the interaction between host and devices, first, the host launches a kernel(s), then the host-launching kernel may or maybe not launch other kernels. A kernel launched from another kernel is called device-launching kernel. These kernels form a kernel call graph, where the start of the graph is the host-launching kernel. The kernel detector only detects host-launching kernels. By monitoring the host code, we can identify kernels launched from the host and subsequently consider them as “used”. This allows us to detect used kernels without the need for GPU-level monitoring.
@@
To launch a kernel from the host, the CUDA driver must call the host function cuModuleGetFunction first. The function takes the name of the kernel to be launched as one of its inputs. It returns a function handle, which is used to execute the kernel. Moreover, cuModuleGetFunction is only called once for each kernel, no matter how many times the kernel is executed, making it ideally suited for our used kernel detection. Inspired by this observation, we implement a hook to the cuModuleGetFunction using Nvidia CUPTI API nvi (d). This hook records the kernel names passed to the function, which are then considered as used kernels. The kernel detector intercepts each host-launching kernel only once and does not intercept device-launching kernels. Consequently, the performance overhead is lower than profiling tools like NSys.
@@
Figure 4 illustrates an example of the kernel detector workflow. When an ML workload invokes the matmul kernel, it first calls the cuModuleGetFunction, which resides in the host code, to launch the kernel. The kernel detector intercepts the call to cuModuleGetFunction and records the kernel name, marked as a used kernel, which is matmul in this example. This kernel may launch other kernels in the device code and form a kernel call graph. However, these device-launching kernels are not detected by the kernel detector.
@@
Figure 4: Structure of device code.
The kernel detector outputs a list of names of host-launching kernels. These kernels are considered as used kernels and are passed to the kernel locator for further processing. Device-launching kernels are also handled by the kernel locator, as we show in the next section.

### v3

@@ Abstract
Software bloat refers to code and features that are not used by a software during runtime. For Machine Learning (ML) systems, bloat is a major contributor to their technical debt, leading to decreased performance and resource wastage. In this work, we present Negativa-ML, a novel tool to identify and remove bloat in ML frameworks by analyzing their shared libraries. Our approach includes novel techniques to detect and locate unnecessary code within GPU code - a key area overlooked by existing research. We evaluate Negativa-ML using four popular ML frameworks across ten workloads over 300 shared libraries. Our results demonstrate that ML frameworks are highly bloated on both the GPU and CPU code side, with GPU code being a primary source of bloat within ML frameworks. On average, Negativa-ML reduces the GPU code size by up to 75% and the CPU code by up to 72%, resulting in total file size reductions of up to 55%. Through debloating, we achieve reductions in peak CPU memory usage, peak GPU memory usage, and execution time by up to 74.6%, 69.6%, and 44.6%, respectively.
@@ 1 Introduction
From personalized recommendations Ko et al. (2022), to healthcare diagnostics Bhardwaj et al. (2017), and autonomous vehicles Parekh et al. (2022), ML is revolutionizing nearly every industrial sector. Besides Large Language Models (LLMs) which enabled natural language understanding and content generation OpenAI Team (2024); Touvron et al. (2023b), smaller models are also now widely deployed in many use-cases, from robots and autonomous vehicles, to cameras and mobile phones. As these models grow in numbers, size, and complexity, managing and optimizing ML systems has become increasingly challenging. LLMs often contain billions of parameters, requiring substantial computational resources and vast amounts of training data. Smaller models on the other hand, are typically deployed in resource-constrained environments. To manage this complexity, substantial technical overhead is introduced to ML systems, exacerbating the issue of software bloat within these systems Zhang et al. (2024).
@@
Software bloat refers to code that is unnecessary for a program during runtime, typically caused by extraneous functions, libraries, or features that do not contribute to the core functionality. Software bloat can cause a range of issues, including decreased performance, increased resource usage, and security vulnerabilities. While bloat can affect any type of software, ML systems have a special capacity for incurring it, as they have all the maintenance problems of traditional code plus an additional set of ML-specific issues, such as boundary erosion, data dependencies, and so on Sculley et al. (2015). As ML models grow in scale and complexity, the bloat in ML systems increases, leading to additional inefficiencies that hinder runtime performance and increase the cost of deployment, particularly in resource-constrained environments Zhang et al. (2024); Jiang et al. (2020).
@@
At the heart of ML systems are ML frameworks, such as TensorFlow Abadi et al. (2016) and PyTorch Paszke et al. (2019). These frameworks provide the essential libraries and tools for model training and inference. ML frameworks are typically developed using multiple programming languages with C++ and CUDA used to implement the core functionalities in order to maximize performance. Both C++ and CUDA code are compiled into shared libraries. Python acts as the frontend, wrapping these core functionalities and enhancing the frameworks’ usability for developers. However, with the ease of use and flexibility that these frameworks provide, they also introduce framework tax Fernandez et al. (2023) - these frameworks can introduce performance overheads and diminish the benefits of new hardware and model architecture advancements. Furthermore, integrating these frameworks also leads to large binary sizes and brings unnecessary overhead for smaller GPUs Jiang et al. (2020).
@@
In this paper, we aim to identify and reduce bloat in ML frameworks by debloating, i.e., removing the bloat, in ML shared libraries. Shared libraries encapsulate the core functionalities of ML frameworks and constitute most of the size of ML framework. These libraries can be hundreds of megabytes to a few gigabytes in size. Debloating ML shared libraries involves the following challenges:
@@
  • ML frameworks rely on some proprietary libraries, such as cuDNN nvi (b) and cuBLAS nvi (c). These libraries are not open-source, making it impossible to perform source code analysis.
@@
  • ML frameworks contain code that runs on both the CPU and GPU. GPU code is overlooked by existing research and no method measure and reduce bloat in it.
@@
  • Although CPU code has a well-defined structure and the bloat in it has been previously studied Brown et al. (2024), the structure of GPU code is not publicly available, making it difficult to analyze and debloat.
@@
To address these challenges, we propose Negativa-ML  [^1] Code will be available at: https://github.com/negativa-ai/negativa-ml, a tool to identify and remove bloat in both CPU and GPU code within ML shared libraries. Leveraging insights into how ML systems execute ML workloads on CPU and GPU, we propose a novel approach to detect the GPU code used by an ML workload with low performance overhead. Subsequently, we locate the file ranges occupied by this GPU code within ML shared libraries. This involves a deep understanding of how GPU code is compiled and how they are organized within a shared library, which is challenging because GPU code structure lacks an open specification. Finally, we utilize a debloating tool that we have previously developed Zhang & Ali-Eldin (2025) to remove bloat from ML shared libraries according to the file ranges located. Our contributions are as follows:
@@
  • We propose an approach to detect GPU code used by ML workloads with low performance overhead.
@@
  • We analyze the structure of GPU code in ML shared libraries and propose a method to locate the file ranges of used GPU code.
@@
  • We extend a debloating tool that we have developed Zhang & Ali-Eldin (2025) to remove bloat in both CPU and GPU code in ML shared libraries.
@@
Our evaluation shows that ML frameworks are highly bloated, with 72% of CPU code and 75% of GPU code being unnecessary for target ML workloads. In addition, 10% of the shared libraries account for up to 90% of the total bloat in the ML frameworks. This bloat not only increases storage overhead but also degrades runtime performance with increased CPU memory usage, GPU memory usage, and longer execution times.
@@ 2.1 ML Frameworks
ML frameworks are software libraries that provide essential tools for building, training, and deploying ML models. Some ML frameworks are general-purpose and designed to support both training and inference for a wide variety of ML models. While others are optimized for specific use cases, such as LLM inference. Two of the most popular general-purpose ML frameworks are TensorFlow Abadi et al. (2016) and PyTorch Paszke et al. (2019), both of which are widely adopted in industry and academia. Although the two frameworks target various ML models, the rise of LLMs has introduced new requirements for ML frameworks, such as efficient KV cache management Kwon et al. (2023); Sinha et al. (2024). LLM inference — which is highly latency-sensitive and resource-demanding — is particularly challenging. To meet these demands, new frameworks specifically designed for LLM inference have been proposed LMDeploy Contributors (2023); Kwon et al. (2023); Wolf et al. (2020); hug .
@@
The standard file format for shared libraries is the Executable and Linkable Format (ELF) ora , which organizes a file into various sections, such as .text and .data. A generic shared library only contains code running on CPU in the .text section. However, ML shared libraries are unique in that they also contain the code designed to run on GPUs. This GPU code is usually included in another section called .nv_fatbin in an ML shared library.
GPU code—also known as device code—significantly increases the size of ML shared libraries compared to traditional shared libraries. The core shared library in PyTorch, torch.so, is 881 MB for the GPU version and 482 MB for the CPU version, nearly double the size. Furthermore, GPU code makes up a substantial portion of these libraries. Analyzing the top four largest shared libraries in PyTorch, GPU code accounts for between 68% to more than 91% of the size of each shared library, as shown in Figure 1. In contrast, CPU code—also known as host code—represents only a small fraction of these libraries. Despite GPU code forming the majority of ML shared libraries, no existing research has investigated bloat within GPU code.
Figure 1: Distribution of CPU code and GPU code in the top 4 largest shared libraries in PyTorch.
@@ 2.2 Software Bloat
Software bloat is mainly caused by unnecessary code in a software, which can be categorized into two types Brown et al. (2024): Type I bloat is universally unnecessary, for example, dead and unreachable code. Type II bloat is end-use dependent, code may be or may not be Type II bloat depending on the use case, for example, unnecessary program features can be used by other programs that use the shared library. Software bloat exists across the entire modern software stack, from operating systems Quach et al. (2017); Kuo et al. (2020), shared libraries Ziegler et al. (2019); Quach et al. (2018); Biswas et al. (2021), to executables Navas & Gehani (2023); Ahmad et al. (2021); Qian et al. (2019) and even containerized applications Rastogi et al. (2017); Zhang et al. (2024). Bloat causes software to grow in size and complexity over time, leading to performance degradation, increased memory consumption, longer startup times, and increased security vulnerabilities, without providing any benefits. Recently, bloat in ML systems has attracted increasing attention. Sculley et al. Sculley et al. (2015) identify that ML systems are easy to accumulate hidden technical debt due to a set of ML-specific issues. Zhang et al. Zhang et al. (2024) show that containerized ML applications are significantly bloated, wasting storage resources and network bandwidth, increasing their attack surface, and slowing down their deployment process.
@@ 2.3 Software Debloating
Software debloating is the process of removing unnecessary code from software to improve efficiency and reduce resource consumption. Many debloating tools have been developed for traditional software, which can be categorized by their debloating targets: source code, binary code, and containerized applications. Source code debloating tools eliminate unused code directly from the source, such as dead or unreachable code, based on specific usage scenarios Brown & Pande (2019); Ye et al. (2021); Azad et al. (2019). Binary debloating tools operate on software binaries, including shared libraries and executables, to remove unnecessary functions or instructions Qian et al. (2019); Ahmad et al. (2021); Ziegler et al. (2019); Agadakos et al. (2019). Container debloating tools target containerized applications, removing unnecessary files in container images Rastogi et al. (2017); Zhang et al. (2023). Most of these tools remove code that is not used by the specific workload, i.e., Type II bloat Qian et al. (2019); Ahmad et al. (2021); Brown & Pande (2019); Rastogi et al. (2017); Azad et al. (2019). While debloating traditional software has been extensively studied, to the best of our knowledge, all existing work focus only on traditional applications where code runs only on CPUs. GPU code, which constitutes a significant portion of ML frameworks, remains unstudied.
One major issue with debloating tools that focus on binary debloating is that they are generally unreliable Brown et al. (2024). To solve their reliability issues, we have recently developed Negativa, a debloating tools that only debloat CPU code Zhang & Ali-Eldin (2025). Negativa demonstrates effectiveness in debloating shared libraries, profiling workloads with the target shared libraries, then removing any code not used by the workload. Negativa’s debloating process involves three phases: detection, location, and compaction. In the detection phase, it identifies CPU functions used by the workload; in the location phase, it locates the used CPU functions within the shared library; and in the compaction phase, it removes unused functions from the shared library and only keeps the used ones.
In this work, we extend Negativa, introducing Negativa-ML, a technique that we have developed to enable GPU code debloating. In addition, we build on Negativa’s capabilities to also debloat CPU code in ML shared libraries. Building on Negativa’s three-phase debloating process, we introduce novel approaches for detecting and locating used GPU code in the detection and location phases. Negativa’s compaction phase is then reused to remove unused GPU code.
@@ 3 Methodology
Figure 2 provides an overview of Negativa-ML. Negativa-ML is composed of two components: the kernel detector and the kernel locator. During the execution of an ML workload, many shared libraries of the ML framework are utilized. The kernel detector monitors the kernel execution of the workload and records the names of the kernels used. The kernel locator analyzes the shared libraries, extracting GPU code within these libraries and finding the file ranges of the used kernels within the shared library. These ranges are subsequently passed to Negativa’s compaction module for debloating. Finally, a debloated ML shared library with reduced CPU code and GPU code is generated.
@@ 3.1 Kernel Detector
The kernel detector is responsible for detecting the kernels used by the ML workload. Although existing tools like Nsight Systems (NSys) nvi (f) can profile GPU code, they are designed primarily to profile and debug GPU performance runtime rather than for kernel detection. These tools, to provide comprehensive information, usually impose a high performance overhead on the target applications by recording data each time a kernel is called. However, for kernel detection, we are only interested in determining whether a kernel has been used, without needing repetitive call data, which results in unnecessary overhead. To offer a more efficient solution, we propose a novel, lightweight approach that specifically detects kernels used within ML shared libraries, achieving low performance overhead. Since kernels execute on GPUs, an intuitive approach would be to monitor GPU execution directly. However, this approach is not feasible because the code execution on GPUs is not directly accessible.
@@
Our approach leverages insights into how ML systems execute ML workloads on both the CPU and the GPU; Considering the interaction between CPU and GPUs, first, the CPU launches a kernel(s), then the CPU-launching kernel may or maybe not launch other kernels. A kernel launched from another kernel is called GPU-launching kernel. These kernels form a kernel call graph, where the start of the graph is the CPU-launching kernel. The kernel detector only detects CPU-launching kernels. By monitoring the CPU code, we can identify kernels launched from the CPU and subsequently consider them as “used”. This allows us to detect used kernels without the need for GPU-level monitoring.
@@
To launch a kernel from the CPU, the CUDA driver must call the CPU function cuModuleGetFunction first. The function takes the name of the kernel to be launched as one of its inputs. It returns a function handle, which is used to execute the kernel. Moreover, cuModuleGetFunction is only called once for each kernel, no matter how many times the kernel is executed, making it ideally suited for our used kernel detection. Inspired by this observation, we implement a hook to the cuModuleGetFunction using Nvidia CUPTI API nvi (d). This hook records the kernel names passed to the function, which are then considered as used kernels. The kernel detector intercepts each CPU-launching kernel only once and does not intercept GPU-launching kernels. Consequently, the performance overhead is lower than profiling tools like NSys.
@@
Figure 4 illustrates an example of the kernel detector workflow. When an ML workload invokes the matmul kernel, it first calls the cuModuleGetFunction, which resides in the CPU code, to launch the kernel. The kernel detector intercepts the call to cuModuleGetFunction and records the kernel name, marked as a used kernel, which is matmul in this example. This kernel may launch other kernels in the GPU code and form a kernel call graph. However, these GPU-launching kernels are not detected by the kernel detector.
@@
Figure 4: Structure of GPU code.
The kernel detector outputs a list of names of CPU-launching kernels. These kernels are considered as used kernels and are passed to the kernel locator for further processing. GPU-launching kernels are also handled by the kernel locator, as we show in the next section.